# Is Quadratic Gravity ghost free because the ghost is only virtual?

### v1

@@ 6.3 Universality properties of the effective action
The above discussion was related to free graviton fields $h_{in\mu\nu}$, which are relevant for studying scattering as they enter in the LSZ formula. It is very difficult to make a discussion about this reduction formula without a proper second quantization of the free or asymptotic fields.
The other important quantity is the partition function $Z(J)$, which is fundamental for calculating the correlation functions in (6.95). The expression consistent with the gauge considered in the original reference [31], described here in (2.24) is
@@
The third massive mode with mass $m_{3}^{2}$ is fundamental for the finding (6.100). The fact that is absent as an external line does not mean that it not contribute to the full correlation functions. In fact, these can be expressed in terms of the effective action $\Gamma_{n}$, whose form does take into account the effect of the third mode.
@@ 7. Final comments and possible future research programs
Before discussing future research programs, some comments about the prescription described between (6.94) and (6.95) are in order. Although the corresponding line of reasoning was described throughout the text, a further justification comes from Chapter 9 of the well known book [155]. In this chapter, these authors develop the path integral formalism starting with creation $a$ and annihilation operators $a^{\dagger}$ and take into account coherent states. On one hand, of the main points of the present work is that the third Stelle mode is gone as an external line. On the other hand, the full path integral $Z(J)$, or the effective action $\Gamma$ depend on the mode, even after the prescription that reference [1]. This may raise a doubt about the consistency of the throwing away procedure of the third mode. More precisely, in Chapter 9, the book [155] develop formulas such as (9-87) or (9.990), which state that the free path integral is
@@
Here $V$ is the interaction term. These authors also derive the scattering LSZ in formula (9-88). These formulas are of course taught in every QFT course. But the particularity of this book presentation is that these formulas are deduced starting with the second quantized theory. In the present context, the interaction $V$ may be interpreted as all the interaction vertices that appear when expanding the Stelle action. The quantity $\Delta(x-y)$ would correspond to the Stelle propagator. This quantity depends on the three mass scales $m_{i}^{2}$ (2.25) with $i=1,2,3$ for each graviton mode. The formalism of [1], which includes the analytic continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ rediscovers, in the euclidean setting, the euclidean version of Stelle gravity. The second quantization presented here is claimed to be in agreement with this continuation, and eliminates the mode with mass scale $m^{2}_{3}$. It may sound odd to work with a theory with a propagator depending on an scale $m_{3}^{2}$ that is not present in the mode expansion. However, this is not an inconsistency, and the comment made in (2.16) is done precisely for clarifying this. The point is that
$\hat{h}_{\mu\nu},\hat{h}_{\alpha\beta}]\neq D_{\alpha\beta,\mu\nu},$
so the left hand side may be independent on $m_{3}^{2}$ while being perfectly possible that the right hand is dependent. It is legal that the propagator depends on a mode which is absent in the free field expansion, although this situation may not be usual in standard Quantum Field Theory.
Note that the LSZ formulas derived in section 4 contains in some intermediary step the commutator of the scalar fields. These steps, adapted to Stelle gravity, will contain the commutator $\hat{h}_{\mu\nu},\hat{h}_{\alpha\beta}]$. However, at the end, these commutators do not appear in the final LSZ formula.
Another issue is that the $S$ matrix presented in this work is unitary. The explanation is simple. In the present work, the Fock states are taken as standard with positive norm, and the usually norm problems are attributed to bad creation and annihilation operators instead to bad norm states. In author´s opinion, the formalism presented here is equivalent to the pioneer ones, in particular [6], [7],[1], [2], [4]-[29]. The use of these states ensures that the $S$ matrix (4.62) is unitary. In fact, consider the Fock spaces $|k_{3}>$ corresponding the third mode and their tensor produces. There appears tensor products with the other components of the Fock space, $|k_{1}>$ and $|k_{2}>$. The Fock space corresponding to the third mode can be considered to decoupled. The unitarity of the $S$ matrix
@@
is not spoiled from this decoupling. There is one state that is decoupled and the other two interact each other.
The only point may be if continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ spoils this property. But this prescription leads to Stelle gravity in an euclidean setting, and one mode is gone as external line. On the other hand, the Fock space is a tensor product, and this procedure, after continuing to Lorenz quantities, leads to a unitary theory with two graviton modes, one massless and one massive.
@@
There are several potential applications related to the present formalism. A first one is related to Bremsstrahlung in gravity theories, and classical limits of gravity. A reference that considered this issue in detail is [82]. The calculation of this effect requires to take the mean value of the wave form $h_{\mu\nu}>$. This calculation, for Quadratic Gravity, involves the prescription presented here, in particular the mapping of the oscillator structure constants $c_{1}=1$, $c_{2}=-1$ to $d_{1}=d_{2}=1$ or, equivalently, $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. There are several works related to this topic, restricted to GR. The work [81] for instance, considers a worldline formalism which may be useful to be generalized to the present context. Other references related to this topic are [51]-[56], [136]-[137]. The works [57]-[60] present interesting results about wave forms, bound states and Hawking radiation which are worthy to be reconsidered in the present context.
@@
The renormalization of the model presented here is related to perturbations along flat spaces. It may seem that applications related to curved backgrounds such as [120]-[121] are outside of the present formalism. However, already in 1973 it was found in reference [75] a connection between quantum Feynman diagrams and the perturbation series of the classical Schwarzschild background. This topic was explored further in [76]-[85], related work is [92]-[98]. These connections between classical and quantum aspects of gravity are worthy to be generalized to the present context. Further applications may be found in [167].
@@ Acknowledgements
The authors are supported by CONICET, Argentina and O. S is supported by the Grant PICT 2020-02181.

### v4

@@ 6.3 Universality properties of the effective action
The above discussion was related to free graviton fields $h_{in\mu\nu}$, which are relevant for studying scattering as they enter in the LSZ formula. The full formula is
$\bigg(\frac{1}{2}+\frac{\kappa^{2}(4\beta-3\alpha)\square}{8}\bigg)P^{0-w}+\frac{\kappa^{4}}{8\lambda}\square[P^{1}-P^{0-w}]\bigg]_{\alpha\beta,\mu\nu}\frac{\delta}{\delta J_{\alpha\beta}(x)}d^{4}x.\bigg\}:Z(J)\bigg|_{J=0},$ (6.83)
with
$h_{in}^{\mu\nu}=\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega(k)}[a_{\mu\nu}a_{p}(k)e^{i\omega(k)t-ik\cdot x}+a_{\mu\nu}^{\ast}a_{p}^{c}(k)e^{-i\omega(k)t+ik\cdot x}]$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{2}(k)}[b_{\mu\nu}b_{p}(k)e^{i\omega_{2}(k)t-ik\cdot x}+b^{\ast}_{\mu\nu}b_{p}^{c}(k)e^{-i\omega_{2}(k)t+ik\cdot x}].$
$\sum_{p}\int\frac{d^{3}k}{(2\pi)^{\frac{3}{2}}\omega_{3}(k)}[c_{\mu\nu}c_{p}(k)e^{i\omega_{3}(k)t-ik\cdot x}+c^{\ast}_{\mu\nu}c_{p}^{c}(k)e^{-i\omega_{3}(k)t+ik\cdot x}].$
The sum is over all the polarizations but the space of physical states $|\Psi>$ eliminates some of them. These states have to be consistent with the Stelle gauge. The algebra of the creation and annihilation is the standard, the wrong (6.81)-(6.82) is changed. The prescription requires to employ the euclidean version of the theory, and then go to the Lorenzian counterpart by continuing to imaginary times.
One may employ other gauges than the Stelle one. The advantage of this gauge is that several universality properties are known for the effective action. The partition function in this gauge is
@@ 7. Final comments and possible future research programs
Before discussing future research programs, some comments about the prescription described between (6.73) and (6.74) are in order. Although the corresponding line of reasoning was described throughout the text, a further justification comes from Chapter 9 of the well known book [159]. In this chapter, these authors develop the path integral formalism starting with creation $a$ and annihilation operators $a^{\dagger}$ and take into account coherent states. The full path integral $Z(J)$, or the effective action $\Gamma$ depend on the mode, even after the prescription that reference [1]. This may raise a doubt about the consistency of the throwing away procedure of the third mode. More precisely, in Chapter 9, the book [159] develop formulas such as (9-87) or (9.990), which state that the free path integral is
@@
Here $V$ is the interaction term. These authors also derive the scattering LSZ in formula (9-88). These formulas are of course taught in every QFT course. But the particularity of this book presentation is that these formulas are deduced starting with the second quantized theory. In the present context, the interaction $V$ may be interpreted as all the interaction vertices that appear when expanding the Stelle action. The quantity $\Delta(x-y)$ would correspond to the Stelle propagator. This quantity depends on the three mass scales $m_{i}^{2}$ (2.22) with $i=1,2,3$ for each graviton mode. The formalism of [1], which includes the analytic continuation $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ rediscovers, in the euclidean setting, the euclidean version of Stelle gravity.
Another issue is that the $S$ matrix presented in this work is unitary. The explanation is that the Fock states are taken as standard with positive norm, and the usually norm problems are attributed to bad creation and annihilation operators instead to bad norm states. In author´s opinion, the formalism presented here is equivalent to the pioneer ones, in particular [6], [7],[1], [2], [4]-[29]. The use of these states ensures that the $S$ matrix (4.43)
@@
is unitary. The evolution of the states is dictated by all the above procedure. The inconsistency is mapped now to the non bounded from below energies or observables mean values. However, the mapping $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$ acting on bra’s $|>$ corrects the signs leading to this unfortunate situation.
@@
There are several potential applications related to the present formalism. A first one is related to Bremsstrahlung in gravity theories, and classical limits of gravity. A reference that considered this issue in detail is [84]. The calculation of this effect requires to take the mean value of the waveform $h_{\mu\nu}>$. This calculation, for Quadratic Gravity, involves the prescription presented here, in particular the mapping of the oscillator structure constants $c_{1}$, $c_{2}$ to $d_{1}$$d_{2}$ or, equivalently, $P_{2}\to iP_{2}$ and $Q_{2}\to iQ_{2}$. There are several works related to this topic, restricted to GR. The work [83] for instance, considers a worldline formalism which may be useful to be generalized to the present context. Other references related to this topic are [53]-[58], [138]-[139]. The works [59]-[62] present interesting results about wave forms, bound states and Hawking radiation which are worthy to be reconsidered in the present context.
@@
The renormalization of the model presented here is related to perturbations along flat spaces. It may seem that applications related to curved backgrounds such as [122]-[123] are outside of the present formalism. However, already in 1973 it was found in reference [77] a connection between quantum Feynman diagrams and the perturbation series of the classical Schwarzschild background. This topic was explored further in [78]-[87], related work is [94]-[100]. These connections between classical and quantum aspects of gravity are worthy to be generalized to the present context.
In addition, the renormalization group of the model, studied with the present techniques may be of clear interest [171]-[173]. Further applications may be found in [174].
@@ Acknowledgements
The author is supported by CONICET, Argentina and by the Grant PICT 2020-02181. I am indebted to Diego Buccio, who pointed out an error in a previous version and whose help was very useful.